OpenStudy (anonymous):

guide me though this please?(: Find the polynomial f(x) that has the roots of -2, 3 of multiplicity 2. Explain how you would verify the zeros of f(x).

OpenStudy (anonymous):

I am really not sure because I am not good with these! @cebroski

OpenStudy (anonymous):

THANK YOU!!

OpenStudy (anonymous):

f(0)=0-1?

OpenStudy (anonymous):

ohhh. okay.. then what?

OpenStudy (anonymous):

x=1

OpenStudy (anonymous):

I have limited time, I need to get to the answer quick! can you explain it quickly all at once? Ill try to read over it afterwards and make myself understand it.

OpenStudy (anonymous):

Find the polynomial f(x) that has the roots of -2, 3 of multiplicity 2. Explain how you would verify the zeros of f(x). f(x) = (x - (-2))*(x - (3))*(x - (3)) f(x) = (x + 2)*(x - 3)*(x - 3) f(x) = x^3 - 4x^2 - 3x + 18 Verify by plugging in the zeros (or roots) in for x and making sure they give you an equation that has both sides equal to each other. Verify that f(-2) = 0 f(3) = 0

OpenStudy (blurbendy):

do you still need help?

OpenStudy (whpalmer4):

After you've got an equation, a further test is to graph it and check the behavior at the roots. A root with an odd multiplicity (1, 3, 5, etc.) will cross the x-axis. A root with an even multiplicity touches the axis and retreats. I see @cebroski has interpreted the question as having only the root 3 have a multiplicity 2. In that case, we'll have one axis crossing and one axis touch.